The sum of three consecutive multiples of 9 is 999.Find these multiples.
Answer by stanbon(75887)

The sum of three consecutive multiples of 9 is 999.Find these multiples.
------
1st: 9(x-1) = 9x-9
2nd: 9x
3rd: 9(x+1) = 9x+9
------
Equation:
sum = 999
---
27x = 999
x = 37
-----
1st: 333-9 = 324
2nd: 333
3rd: 333+9 = 342
========================
Cheers,
Stan H.
=======================